OpenStudy (anonymous):

trig help.. solve each equation for solutions in the interval [0, 360) tan θ +1 = √3 + √3cotθ? any help would be appreciated!

OpenStudy (jdoe0001):

$$ tan(\theta)+1 = \sqrt{3}+\sqrt{3}cot(\theta)\\ tan(\theta)+1 = \sqrt{3}(1+cot(\theta))\\ \cfrac{tan(\theta)+1}{1+cot(\theta)} = \sqrt{3} $$

OpenStudy (jdoe0001):

so, from the trig identities tan = sin/cos cot = cos/sin we'll expand that to

OpenStudy (jdoe0001):

$$ \cfrac{ \cfrac{sin(\theta)}{cos(\theta)}+1} { 1+\cfrac{cos(\theta)}{sin(\theta)}}\\ \implies \cfrac{ \cfrac{\color{blue}{sin(\theta)+cos(\theta)}} {cos(\theta)}} { \cfrac{\color{blue}{sin(\theta)+cos(\theta)}} {sin(\theta)}}\\ $$

OpenStudy (jdoe0001):

well all that \(\huge = \sqrt{3}\)

OpenStudy (jdoe0001):

see any like terms to cancel out?

OpenStudy (anonymous):

yes, but where did the +1's go?

OpenStudy (jdoe0001):

in the fraction addition s/c +1 GCF of both fractions, is c so (s+c)/c

OpenStudy (jdoe0001):

you're really adding sine/cosine + 1/1

OpenStudy (anonymous):

okay, thanks! i get it now :)

OpenStudy (jdoe0001):

$$ \cfrac{ \cfrac{\color{blue}{sin(\theta)+cos(\theta)}} {cos(\theta)}} { \cfrac{\color{blue}{sin(\theta)+cos(\theta)}} {sin(\theta)}}= \sqrt{3}\\ \implies tan(\theta) = \sqrt{3} $$

OpenStudy (jdoe0001):

now you just need to find in your Unit Circle, what angle has a tangent of \(\sqrt{3}\) or just arctan both sides

OpenStudy (anonymous):

okay, thanks for your time :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

@jdoe0001 so the answer would be π/3 and 4π/3 right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

I and II quadrants, only place where tangent is positive

OpenStudy (jdoe0001):

woops, I and III rather

OpenStudy (anonymous):

that's what i got thanks again!

OpenStudy (jdoe0001):

yw